3. A ball is thrown vertically upward with a speed of 30 ft/s from the edge of a cliff 50 ft above sea level with what velocity with it hit the water?





a. 32.2 ft/s b. 58.6 ft/s c. 64.2 ft/s d. 52.3 ft/s







1
Expert's answer
2022-02-23T08:31:08-0500

The law of conservation of energy says

"\frac{mv_i^2}{2}+mgh=\frac{mv_f^2}{2}"

Hence, the final velocity

"v_f=\sqrt{v_i^2+2gh}\\=\sqrt{30^2+2*32.17*50}=64.2\:\rm ft/s"

Answer: c. 64.2 ft/s